# MIAT: Maneuver-Intention-Aware Transformer for Spatio-Temporal Trajectory Prediction

### v1

@@ VI Conclusion and Future Work
We introduce Maneuver-Intention-Aware Transformer (MIAT) for vehicle trajectory prediction. MIAT employs parallel self-attention to model long-range socio-temporal dependencies while explicitly accounting for driving behaviors through maneuver-classification. Experiments on real-world dataset confirm that MIAT not only achieves strong short-horizon accuracy, but also markedly outperforms the baselines at longer horizons. Moreover, the Transformer’s parallelism provides computational benefits, making MIAT feasible for real-time traffic engineering applications. By unifying trajectory and maneuver objectives in a single loss function, MIAT robustly balances near-term precision and extended-horizon reliability.
There exist many future research directions. First, we are interested in testing our approach in more complex and larger scenarios [30, 31] that can be calibrated using mobile data to reflect real-world traffic conditions [32, 33, 34]. Second, we would like to incorporate additional generic information such as traffic state predictions [35, 36] and vehicle trajectories [37] into our framework for potential improvement. Third, we want to improve the resilience of our approach by considering adversarial attacks [38, 39, 40] into account. Finally, we are interested in studying the integration of our approach with mixed traffic control [41, 42, 43, 44].
Acknowledgments
This research is supported by NSF Award IIS-2153426. The authors would like to thank Bibek Poudel for his efforts and NVIDIA and the Tickle College of Engineering at UTK for their support.

### v3

@@ VI Conclusion and Future Work
This work introduces two key advancements over prior intention-aware models. First, a tunable intention awareness control mechanism is presented. Unlike approaches that use a fixed weight loss for trajectory and maneuver objectives, this mechanism allows for a dynamic balance between trajectory loss and maneuver loss, effectively balancing short-term precision and long-term consistency. Notably, a deliberate 200 times scale of the maneuver loss yielded an 11.1% better performance in 5-second horizon compared to STDAN, demonstrating a promising method for enhancing long-horizon forecasting.
Second, a Transformer architecture was integrated, replacing conventional LSTM backbones. This architectural shift enables MIAT to effectively model long-range spatio-temporal dependencies, providing a robust foundation for an intention-aware learning strategy. This approach is crucial for achieving reliable performance over extended horizons, directly addressing the inherent limitations of traditional sequential models. Ultimately, MIAT offers a robust framework and competitive methodology for real-time trajectory prediction systems, particularly in applications where long-horizon reliability is paramount.
Future work involves enhancing MIAT by incorporating graph-based reasoning to leverage road topology. The framework will also be extended to generate multi-modal trajectory distributions to capture a wider range of plausible outcomes, and its robustness will be validated on larger, more complex datasets.
Acknowledgements
This research is supported by NSF IIS-2153426. The authors thank NVIDIA and the Tickle College of Engineering at University of Tennessee, Knoxville for their support.